OpenStudy (babyslapmafro):

Please help me find the interval of convergence of the given power series. (click to see)

OpenStudy (babyslapmafro):

\[\sum_{n=1}^{\infty} \frac{ (-1)^{n} }{ n }x ^{n}\]

OpenStudy (babyslapmafro):

@abb0t

OpenStudy (anonymous):

Use the ratio test: \[\lim_{n\to\infty}\left|\frac{(-1)^{n+1}x^{n+1}}{n+1}\cdot\frac{n}{(-1)^nx^n}\right|=\lim_{n\to\infty}\frac{n}{n+1}|x|\] A series converges by the ratio test if this limit is less than 1.